OpenStudy (anonymous):

how do you graph y=3=2(x-1)

OpenStudy (anonymous):

that's not a valid equation. you should just have one equals sign ... like maybe y = 3 + 2(x-1)

OpenStudy (anonymous):

argumentative: if y=3 and 2(x-1)=3 then y=3=2(x-1), but just ignore me! it does look like maybe it should be y=3-2(x-1) or something.

OpenStudy (anonymous):

idk thats what it said lol